OpenStudy (anonymous):

Find the sum of a finite geometric sequence from n = 1 to n = 8, using the expression −2(3)^n − 1. 1,083 −6,560 −4,374 6,559

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

sum of finite geometric sequence : \(\huge \frac{a(r^n-1)}{r-1} \)

ganeshie8 (ganeshie8):

start by finding the first term \(a\) and the common ratio \(r\) by interpreting given expression :- \( −2(3)^{n − 1}\)

ganeshie8 (ganeshie8):

compare it wid the general term :- \(ar^{n-1}\) a = ? r = ?

ganeshie8 (ganeshie8):

a = -2, r = 3 right ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

oh ok

ganeshie8 (ganeshie8):

we just compared :) plug them in the sum formula

ganeshie8 (ganeshie8):

\(\huge \frac{a(r^n-1)}{r-1}\) a = -2 r = 3 n = 8 plug and feed it to ur calculator

OpenStudy (anonymous):

-13120/2=-65560

OpenStudy (anonymous):

right? @ganeshie8

ganeshie8 (ganeshie8):

if u mean -6560 then yes :)

OpenStudy (anonymous):

Thank You

ganeshie8 (ganeshie8):

np :)